A basketball player threw the ball with an initial velocity of 30 m/s 30°above the horizontal.




a.How long it took the ball to land?



b.How high the ball flew?



c.What was its range?

1
Expert's answer
2021-12-03T12:43:40-0500

given properties:

"v_o=30 \frac{m}{s}"

"\alpha = 30^o"

a) "t = \large\frac{2v_osin\alpha}{g}" "=\large\frac{2*30*0.5}{10}" "=3s"

b) "H_m = \large\frac{v_o^2sin^2\alpha}{2g}" "= \large\frac{900*0.25}{2*10}" "=11.25m"

c) "L=\large\frac{v_o^2sin2\alpha}{g}" "=\large\frac{900*0.86}{10}" "=78m"